# Understanding In-context Learning of Addition via Activation Subspaces

### v1

@@ Abstract
footnotetext: Equal advising.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable in-context learning (ICL) capabilities, where they can adapt to various new tasks through few-shot demonstrations in the prompt. Motivated by this, researchers have constructed detailed models of in-context learning for small synthetic language models (Garg et al., 2022; von Oswald et al., 2023; Bai et al., 2023; Zhang et al., 2023; Nichani et al., 2024) as well as coarser-grained treatments of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024). However, little is known about the fine-grained computational structure of in-context learning for large models.
@@
To understand the mechanism behind in-context add-k, we study it in Llama-3-8B. We first discover that only a small number of model components are responsible for add-k: out of 1024 attention heads, only three do most of the work (§3). We find the task information for each head (i.e., the number $k$) is encoded in a six-dimensional subspace; zooming into one of the heads, we further decompose it into a four-dimensional space tracking the unit digit using sine and cosine functions and a two-dimensional one tracking the overall magnitude (§4). We also examine how this head aggregates information from individual in-context examples (§5).
@@
Since these three heads are responsible for computing add-k, they must extract signal about $k$ from individual examples and then represent $k$ in the residual stream. To understand how $k$ is represented, we first find that the effect of each head factors through only a six-dimensional subspace (out of $128$ in total) via principal component analysis (PCA) (§4.1). Furthermore, zooming into the six-dimensional subspace of one head, we find that it factors into a four-dimensional subspace tracking the unit digit through trigonometric functions (with frequencies $2,5,10$), and a two-dimensional subspace tracking the coarse-grained magnitude (the ten digit) (§4.2-4.3). This offers insight on how LLMs internally represent addition, aligning with recent findings showing that (non-ICL) addition is encoded via trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025).
Finally, we examine how this head aggregates information from individual in-context examples. We find that the information is mostly extracted from the $y_{i}$ token positions and that this extracted signal encodes a vector representation of $y_{i}-x_{i}$5.2). Moreover, we identify a self-correction mechanism: if the signal extracted from one token has an error, later tokens are likely to write in the opposite direction of the error (§5.3). This means that ICL goes beyond simple averaging of inputs and has stateful dynamics.
@@ 2.1 Model and Task
We begin by specifying the model and task studied in this paper.
@@
For a constant $k$, the add-k task is to add $k$ to a given input integer $x$ to predict $y=x+k$. In an $n$-shot ICL prompt, we provide $n$ examples of the format “$x_{i}\rightarrow y_{i}$" (where each $y_{i}=x_{i}+k$). The examples are concatenated in the prompt using the separator “##\##", followed by a query “$x_{q}\rightarrow$”; see Figure 1 for an example of a five-shot ICL prompt. A key advantage of this family of tasks is that ICL prompts for different tasks only differ in $k$ (i.e. $y_{i}-x_{i}$ in the prompts) but not the input domain, in contrast to most previous studies (Todd et al., 2024; Hendel et al., 2023).
@@
    zero-shot prompt, where there are no demonstration examples: the prompt is “$x_{q}\rightarrow$" for some $x_{q}$.
@@ 3.2 Further Refinement via Ablations
Focusing on two layers via layer-wise ablation. To efficiently narrow down the important heads, we first perform ablations at the level of layers. From Figure 3(a), we observe that the significant heads $\mathcal{H}_{\text{sig}}$ are concentrated primarily in the middle and late layers. We speculate that heads in the late layers mainly contribute to formatting the output, as they appear too late in the computation to meaningfully interact with the query. After trying different sets of layers, we found that ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while ablating any other combination of layers causes negligible drops in accuracy (Table 3). After these ablations, only $11$ heads located in layers $13$ and $15$ remain.
@@
$(15,2)\text{:=Head1}$ $6$   0.85
$(15,1)\text{:=Head2}$ $5$   0.83
$(13,6)\text{:=Head3}$ $5$   0.66
@@ 4 Characterizing the Aggregator Subspace
For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL examples. Our next goal is to understand how task information is represented in the activation space. Since we have identified three aggregator heads that carry almost all of this information, we can now focus on analyzing the representation space of these three heads.
@@ 4.1 Localizing to Six-dimensional Subspace
To reduce the $128$-dimensional head activation to a more tractable space to study, we first perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance (Figure 4). We then check that the function vectors found earlier remain effective after projecting onto the subspace. Specifically, we replace each head vector $h_{k}$ with its projection onto the subspace $\tilde{h}_{k}$ to obtain a new function vector $\tilde{v}_{k}=\sum_{h\in\mathcal{H}}\tilde{h}_{k}$ (a variant of $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ in §2.2). We find that $\tilde{v}_{k}$ has intervention accuracy $0.76$, which is close to the intervention accuracy of $0.79$ before projection (§3.2). Thus, we confine our study to this concise six-dimensional subspace for each head. The rest of the section focuses on head 1 for simplicity.
@@ 4.3 Decomposing to Subspaces Encoding Subsignals
By construction, each of the feature directions (viewed as linear projections of $h_{k}$) are periodic functions of $k$. Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo $T$”. Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
We verify these hypotheses through causal intervention. In Figure 7, we validate hypothesis (ii): we establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal. We include analogous experimental results for the two other hypotheses in Appendix B.
Figure 7: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto the “unit subspace” (left) and out of the “unit subspace” (right). Projecting onto the unit subspace results in a low unit-digit error rate even when final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully accounts for the final-answer errors. This confirms that the unit subspace specifically encodes the unit-digit signal.
@@ 5 Signal Extractors of ICL Examples
In this section, we find that: (1) the signal is primarily gathered from the $y_{i}$ tokens in each example $x_{i}\rightarrow y_{i}$; (2) each example $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context examples with conflicting information; and (3) when all examples $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior. For simplicity, we focus on head 1 in this section.
@@ 5.2 Signal Extractor for each Example
We next examine what specific information is extracted from each of these tokens. To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},h_{k}\rangle$ between the extracted information and the head vector for each task $k$. In Figure 9, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$. We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
Figure 9: For a random mixed-$k$ ICL prompt, the inner product between the projected signal from each $y_{i}$ and each head vector $h_{k}$ (projected onto the subspace and normalized to have unit norm). For each example $x_{i}\rightarrow y_{i}$, the inner product peaks for $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
@@ 5.3 Signal Correlation among Examples
Statistics over tasks Sum of $\mathbf{negative}$ correlations Sum of $\mathbf{positive}$ correlation
@@ 6 Related Work
Our work builds on a growing body of research that aims to uncover circuits and internal computations of language models (Elhage et al., 2021; Wu et al., 2023; Wang et al., 2022; Feng & Steinhardt, 2023; Gould et al., 2023).
@@
A recent line of work examines how LLMs perform arithmetic (Stolfo et al., 2023; Nikankin et al., 2024), and in particular addition (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024; Kantamneni & Tegmark, 2025). Zhou et al. (2024) find that pre-trained LLMs perform addition using Fourier features, and Kantamneni & Tegmark (2025) find that mid-sized LLMs compute addition using a “clock” algorithm via a helix representation of numbers. Different from them, our work is the first to analyze addition for in-context learning in a large pretrained model, and we find similar representation structures to Zhou et al. (2024) and Kantamneni & Tegmark (2025).
@@
Researchers have constructed detailed models of in-context learning (ICL) for small transformer models (Garg et al., 2022; von Oswald et al., 2023; Bai et al., 2023; Zhang et al., 2023; Nichani et al., 2024), as well as coarser-grained treatments for large pretrained models, attributing ICL performance to either induction heads (Olsson et al., 2022; Singh et al., 2024; Crosbie & Shutova, 2025; Bansal et al., 2023) or function vector (FV) heads (Todd et al., 2024; Hendel et al., 2023). Yin & Steinhardt (2025) compares the two types of heads and finds that few-shot ICL performance depends primarily on FV heads. Motivated by this, we perform a detailed analysis on the mechanisms behind a family of ICL addition tasks through a novel optimization method inspired by function vectors, which achieves better performance than the method in Todd et al. (2024). Unlike the tasks studied in Todd et al. (2024), our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information.
@@
Our methodological approach draws on recent developments in the causal analysis of neural networks that aims to understand the causal influence of model components to its behavior by performing targeted interventions on the internal states of models (Geiger et al., 2021; Vig et al., 2020). In particular, our methods draw inspiration from the causal mediation analysis used in Todd et al. (2024), activation patching (Meng et al., 2022), and causal scrubbing (Chan et al., 2022).
@@ Acknowledgments
Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers. We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.8673$, while other combinations lead to negligible drops in accuracy, especially when mean-ablating the layer $13$ and $15$ (in blue).
Appendix B Additional figures in §4.3
Figure 10: Head (15,2) projected onto parity direction
Figure 11: Head (15,2) projected out of parity direction
Figure 12: Head (15,2) projected onto mag subspace
Figure 13: Head (15,2) projected out of mag subspace

### v2

@@ Abstract
footnotetext: Equal advising.\$*\$\$*\$footnotetext: Correspondence to: Xinyan Hu <xinyanhu@berkeley.edu>.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable in-context learning (ICL) capabilities, where they can adapt to various new tasks through few-shot demonstrations in the prompt. Motivated by this, researchers have constructed detailed models of in-context learning for small synthetic language models (Garg et al., 2022; Akyürek et al., 2023; von Oswald et al., 2023) as well as coarser-grained treatments of large pretrained models (Olsson et al., 2022; Hendel et al., 2023; Todd et al., 2024). However, little is known about the fine-grained computational structure of in-context learning for large models.
@@
To understand the mechanism behind in-context add-k, we study it in Llama-3-8B. We first discover that only a small number of model components are responsible for add-k: out of 1024 attention heads, only three do most of the work (§3). We find the task information (i.e., the number $k$) for each head is encoded in a six-dimensional subspace; zooming into each head, we further decompose it into a four-dimensional space tracking the unit digit using sine and cosine functions and a two-dimensional one tracking the overall magnitude (§4). We also examine how this head aggregates information from individual in-context examples (§5).
@@
Since these three heads are responsible for computing add-k, they must extract signal about $k$ from individual examples and then represent $k$ in the residual stream. To understand how $k$ is represented, zooming into each head, we first find that the effect of each head factors through only a six-dimensional subspace (out of $128$ in total) via principal component analysis (PCA) (§4.1). Furthermore, we find that it factors into a four-dimensional subspace tracking the unit digit through trigonometric functions, and a two-dimensional subspace tracking the coarse-grained magnitude (§4.2-4.3). This offers insight on how LLMs internally represent addition, aligning with recent findings showing that (non-ICL) addition is encoded via trigonometric (Zhou et al., 2024) or helical functions (Kantamneni & Tegmark, 2025).
Finally, we examine how this head aggregates information from individual in-context examples. We find that the information is mostly extracted from the $y_{i}$ token positions, and that this extracted signal encodes a vector representation of $y_{i}-x_{i}$5.2). Moreover, we identify a self-correction mechanism: if the signal extracted from one token has an error, later tokens are likely to write in the opposite direction of the error (§5.3). This means that ICL goes beyond simple averaging of inputs and has stateful dynamics.
@@ 2.1 Model and Task
We begin by specifying the model and the task studied in this paper.
@@
For a constant $k$, the add-k task is to add $k$ to a given input integer $x$ to predict $y=x+k$. In an $n$-shot ICL prompt, we provide $n$ examples of the format “$x_{i}\rightarrow y_{i}$” (where each $y_{i}=x_{i}+k$). The examples are concatenated in the prompt using the separator “##\##”, followed by a query “$x_{q}\rightarrow$”; see Figure 1 for an example of a five-shot ICL prompt. A key advantage of this family of tasks is that ICL prompts for different tasks only differ in $k$ (i.e. $y_{i}-x_{i}$ in the prompts) but not the input domain, in contrast to most previous studies (Todd et al., 2024; Hendel et al., 2023).
@@
    zero-shot prompt, where there are no demonstration examples: the prompt is “$x_{q}\rightarrow$” for some $x_{q}$.
@@ 3.2 Further Refinement via Ablations
Focusing on two layers via layer-wise ablation. To efficiently narrow down the important heads, we first perform mean-ablations at the level of layers. From Figure 3(a), we observe that the significant heads $\mathcal{H}_{\text{sig}}$ are concentrated primarily in the middle and late layers. We speculate that heads in the late layers mainly contribute to formatting the output, as they appear too late in the computation to meaningfully interact with the query. After trying different sets of layers, we found that mean-ablating all significant heads outside layers $13$ and $15$ still achieves an intervention accuracy of $0.83$, while mean-ablating any other combination of layers causes negligible drops in accuracy (Appendix A, Table 3). After these ablations, only $11$ heads located in layers $13$ and $15$ remain.
@@
[…]
@@ 4 Characterizing the Aggregator Subspace
For the model to perform add-k, it has to infer the task information (the number $k$) from the ICL examples. Our next goal is thus to understand how task information is represented in the activation space. Since we have identified three aggregator heads that carry almost all of this information, we can now focus on analyzing the representation space of these three heads.
@@ 4.1 Localizing to Six-dimensional Subspace
To reduce the $128$-dimensional head activation to a more tractable space to study, we first perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance (Figure 4). We then check that the function vectors found earlier remain effective after projecting onto the subspace. Specifically, we replace each head vector $h_{k}$ with its projection onto the subspace $\tilde{h}_{k}$ to obtain a new function vector $\tilde{v}_{k}=\sum_{h\in\mathcal{H}}\tilde{h}_{k}$ (a variant of $v_{k}=\sum_{h\in\mathcal{H}}h_{k}$ in §2.2). We find that $\tilde{v}_{k}$ has intervention accuracy $0.76$, which is close to the intervention accuracy of $0.79$ before projection (§3.2). Thus, we confine our study to this concise six-dimensional subspace for each head. The rest of the section focuses on head 1 and we defer analogous results for head 2 and 3 to appendix B.
@@ 4.3 Decomposing to Subspaces Encoding Subsignals
By construction, the coordinate function of each feature directions (viewed as linear projections of $h_{k}$) is a periodic function of $k$. Mathematically, a feature direction with period $T$ carries task information from the head vectors with “modulo $T$”. Based on this, we hypothesize: (i) the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task; (ii) the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$; (iii) the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
We verify these hypotheses through causal intervention. In Figure 7, we validate hypothesis (ii): we establish (1) sufficiency by showing that projecting a head vector onto the subspace preserves the relevant task signal; and (2) necessity by showing that projecting a head vector out of the subspace (i.e., onto its orthogonal complement) destroys the relevant task signal. We include analogous experimental results for the two other hypotheses of head 1 in Appendix B.2.
Figure 7: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto the “unit subspace” (left) and out of the “unit subspace” (right). Projecting onto the unit subspace results in a low unit-digit error rate even when final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully account for the final-answer errors. This confirms that the unit subspace specifically encodes the unit-digit signal.
@@ 5 Signal Extractors of ICL Examples
In this section, we find that: (1) the signal is primarily gathered from the $y_{i}$ tokens in each example $x_{i}\rightarrow y_{i}$; (2) each example $x_{i}\rightarrow y_{i}$ individually contributes a signal $y_{i}-x_{i}$ in the subspace even on “mixed” in-context examples with conflicting task information; and (3) when all examples $x_{i}\rightarrow y_{i}$ share the same value for $y_{i}-x_{i}$, the extracted signals exhibit a self-correction behavior. For simplicity, we focus on head 1 in this section.
@@ 5.2 Signal Extractor for each Example
We next examine what specific information is extracted from each of these tokens. To do so, we measure the inner product $\langle W_{h}O_{h}V_{h}z_{t},\tilde{h}_{k}\rangle$ between the extracted information and the head vector (projected onto the subspace and normalized to have unit norm) for each task $k$. In Figure 9, we plot this quantity for a random mixed-$k$ ICL prompt for each token $y_{i}$ ($i\in\{1,\ldots,5\}$) and each task $k\in\{1,\ldots,30\}$. We find that the inner product consistently peaks at $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
Figure 9: For a random mixed-$k$ ICL prompt, the inner product between the projected signal from each $y_{i}$ and each head vector (projected onto the subspace and normalized to have unit norm) $\tilde{h}_{k}$. For each example $x_{i}\rightarrow y_{i}$, the inner product peaks for $k=y_{i}-x_{i}$, indicating that the model extracts the information of $y_{i}-x_{i}$ from the corresponding example $x_{i}\rightarrow y_{i}$.
@@ 5.3 Signal Correlation among Examples
Statistics over tasks Sum of $\mathbf{negative}$ correlations Sum of $\mathbf{positive}$ correlations
@@ 6 Related Work
Our work builds on a growing body of research that aims to uncover circuits and internal computations of language models (Olah et al., 2020; Wang et al., 2022; Wu et al., 2023; Cunningham et al., 2023; Feng & Steinhardt, 2023; Gould et al., 2023; Park et al., 2024). Below we discuss three particular threads that are most relevant to this paper.
@@
A recent line of work examines how LLMs perform arithmetic (Stolfo et al., 2023; Hanna et al., 2023; Nikankin et al., 2024; Maltoni & Ferrara, 2024), and in particular addition (Nanda et al., 2023; Zhong et al., 2023; Zhou et al., 2024; Kantamneni & Tegmark, 2025). Zhou et al. (2024) find that pre-trained LLMs perform addition using Fourier features, and Kantamneni & Tegmark (2025) find that mid-sized LLMs compute addition using a “clock” algorithm via a helix representation of numbers. Different from them, our work is the first to analyze addition for in-context learning in LLM, and interestingly we find similar representation structures to Zhou et al. (2024) and Kantamneni & Tegmark (2025).
@@
Researchers have constructed detailed models of in-context learning (ICL) for small transformer models in standard supervised learning problems such as linear regression (Garg et al., 2022; Akyürek et al., 2023; Zhang et al., 2023; Li et al., 2023; Wu et al., 2024), as well as more complex settings (von Oswald et al., 2023; Bai et al., 2023; Bietti et al., 2023; Reddy, 2023; Guo et al., 2023; Nichani et al., 2024). For large pretrained models, there exist coarser-grained treatments attributing ICL performance to either induction heads (Olsson et al., 2022; Singh et al., 2024; Crosbie & Shutova, 2025; Bansal et al., 2023) or function vector (FV) heads (Todd et al., 2024; Hendel et al., 2023). Yin & Steinhardt (2025) compares the two types of heads and finds that few-shot ICL performance depends primarily on FV heads. Motivated by this, we study function vector heads in details for a family of ICL tasks, introducing a novel optimization method, which achieves better performance than the method in Todd et al. (2024). Another difference from Todd et al. (2024) is that our tasks have the same input domain, ensuring that the ICL prompts for different tasks differ only in the task information.
@@
There has been a line of research that proposes methods to understand the causal influence of model components on model behavior, such as by probing (Conneau et al., 2018; Hewitt & Manning, 2019; Clark et al., 2019). Our methodological approach follows recent developments in revealing causal effects of model components by interventions on internal states of models (Vig et al., 2020; Geiger et al., 2021). In particular, we draw inspiration from causal mediation analysis used in Todd et al. (2024), activation patching (Meng et al., 2022), and causal scrubbing (Chan et al., 2022).
@@ Acknowledgments
In §3.2, we did systematic ablation studies to narrow down the tasks to three main heads from $33$ significant heads. The first step of the ablation studies is layer-wise ablation, where we mean-ablate the significant heads in a subset of layers. We include the experimental results here, which narrow down to layer 13 and 15.
@@
Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers. We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.87$, while other combinations lead to substantial drops in accuracy, especially when mean-ablating layers $13$ and $15$ (in blue).
Appendix B Additional figures in §4
In §4.2 and §4.3, we focused on head 1 for simplicity of exposition. Here we include the analogous experimental results for head 2 and 3.
B.1 Additional figures for §4.2
In §4.2, we found six trigonometric functions that can be linearly fitted by the coordinate functions of head 1 (Figure 6). Here we show analogous experimental results for head 2 and 3.
Figure 10: The coordinate functions (inner product with feature directions) of head 2’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
Figure 11: The coordinate functions (inner product with feature directions) of head 3’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
B.2 Additional figures for §4.3
In §4.3, we showed experimental results validating the hypothesis (ii) for head 1 (Figure 7). Here we show analogous experimental results validating the hypotheses (i) and (iii) for head 1, as well as all three hypotheses for heads 2 and 3. Recall that the three hypotheses are as follows:
  1. (i)
    the feature direction corresponding to period two, which we call the “parity direction”, encodes the parity of $k$ in the add-k task;
  2. (ii)
    the subspace spanned by the feature directions with periods $2,5,10$, which we call the “unit subspace”, encodes the unit digit of $k$;
  3. (iii)
    the subspace spanned by by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
We first show the experimental results validating hypotheses (i) and (iii) for head 1. Projecting out of the parity direction doesn’t lead to high errors for the parity and the final answer across all tasks, which might be because parity is relatively easy to obtain (e.g., random choice leads to 0.5 accuracy).
(a) Head 1 projected onto parity direction
(b) Head 1 projected out of parity direction
Figure 12: Validation of hypothesis (i) for head 1.
(a) Head 1 projected onto magnitude subspace
(b) Head 1 projected out of magnitude subspace
Figure 13: Validation of hypothesis (iii) for head 1.
We then show analogous results for head 2.
(a) Head 2 projected onto parity direction
(b) Head 2 projected out of parity direction
Figure 14: Validation of hypothesis (i) for head 2.
(a) Head 2 projected onto unit subspace
(b) Head 2 projected out of unit subspace
Figure 15: Validation of hypothesis (ii) for head 2.
(a) Head 2 projected onto magnitude subspace
(b) Head 2 projected out of magnitude subspace
Figure 16: Validation of hypothesis (iii) for head 2.
Finally, we show analogous results for head 3. The evidence for head 3 is slightly weaker than the first two heads, which aligns with the fact that the intervened accuracy of head 3 is slightly weaker than the ones of the first two heads (Tabel 1).
(a) Head 3 projected onto parity direction
(b) Head 3 projected out of parity direction
Figure 17: Validation of hypothesis (i) for head 3.
(a) Head 3 projected onto unit subspace
(b) Head 3 projected out of unit subspace
Figure 18: Validation of hypothesis (ii) for head 3.
(a) Head 3 projected onto magnitude subspace
(b) Head 3 projected out of magnitude subspace
Figure 19: Validation of hypothesis (iii) for head 3.